Equilbrium constant Kc fr the reaction:


So2(g) + NO2(g) -> SO3(g) + NO(g) is 16. If 1 mole each of all 4 gases is taken in 1dm3 vessel , equil. conc. of NO would be?

Grade:11

3 Answers

APURV GOEL
12 years ago

Take moles of So2 and No2 as 1-x and that of So3 and No as 1+x.

equate the kc and get the value of x and you will get the moles of No by 1+x and if volume is 1 then 1+x will be its conc.

we get

(1+x)^2/(1-x)^2 = 16

solve to get x = 0.6

and the ans will be 1.6.

APURV GOEL
12 years ago

Take conc of So2 and No2 as 1-x and conc of So3 and No as 1+x and equate it to kc

as (1-x/1+x) = 16

and get x = 0.6

and conc of No at eq. will be 1.6

manish nayak
12 years ago

u can assume tat x moles hv rktd

x2/(1-x)2=16

so ans is 0.8